Consider that you have a balloon containing 2.00 moles of CO and 1.00 mole of O2 which is in a room that has a temperature of 26.6oC and a pressure of 1.00 atm. Then, the following reaction occurs inside the balloon to completion: 2 CO(g) + O2(g) -> 2 CO2(g). Calculate the change in work due to the reaction occurring inside the balloon.

Enter your answer to three significant figures and in units of kJ.

-Enter 3 significant figures unless otherwise told.

-Use scientific notation if necessary using the letter "e" followed by the power (ex: 1e-9)

-Use positive or negative signs for energy when necessary

-Use the following conversion if needed: 1 L-atm = 101.3 J
Answer #1

2 CO(g) + O2(g) -----------> 2 CO2(g)

temperature = 26.6 + 273.15 = 299.75 K

delta n = 2 - (2 + 1) = -1

w = delta n x R T

   = - 1 x 8.314 x 299.75

= 2492.12 J

change in work = 2.49 kJ